### v1

# Cascade Speculative Drafting for Even Faster LLM Inference

Abstract
Speculative decoding enhances the efficiency of large language models (LLMs) by leveraging a draft model to draft for a larger target model to review. However, drafting in speculative decoding involves slow autoregressive generation and generating tokens of different importance with the same time allocation. These two inefficiencies lead to its suboptimal performance. To address this issue, we introduce Cascade Speculative Drafting (CS. Drafting), a novel approach that employs two types of cascades. The Vertical Cascade eliminates autoregressive generation from neural models. The Horizontal Cascade constitutes efficient time allocation in drafting with its optimality supported by our theoretical analysis. Combining both cascades, our CS. Drafting algorithm has achieved up to 72 percent additional speedup over speculative decoding in our experiments while keeping the same output distribution.
Machine Learning, ICML
[…]
1 Introduction
The advent of Large Language Models (LLMs) like GPT-4 (OpenAI, 2023) has marked a significant milestone in the field of natural language processing (NLP). These models have not only excelled in various NLP tasks but also found widespread application in user-interactive settings such as chatbots and virtual assistants. This application involves an extremely high number of users of up to hundreds of millions daily. To service in real-time at this scale, a low latency system is not only cost-saving but also crucial for keeping the service running. In addition, the sheer scale of the service making even a slight improvement over the latency of LLMs can contribute greatly to the community. Consequently, optimizing LLM latency has become a crucial and impactful research focus.
Unfortunately, LLMs currently suffer from high latency especially for long-form generation as LLM can only generate one token at each run. The challenge is especially great during applications such as chatbots or virtual assistants due to the long generation of hundreds or even thousands of tokens. An emerging solution, known as speculative decoding (Leviathan et al., 2023; Chen et al., 2023), shows potential to mitigate this issue. In speculative decoding, a draft model (which is smaller and faster) generates $k$ tokens in each step(with $k$ being a hyperparameter), and these tokens are then reviewed by a target model (which is larger and slower) in parallel. The target model will accept any tokens aligned with its output and further generate one token; the process happens all in a single run while maintaining the output distribution identical to the autoregressive one by the target model itself. As a larger number of expected accepted tokens per run indicates fewer target model runs, the efficiency of speculative decoding can be attributed to two factors: the expected number of tokens accepted by the target model per run, and the autoregressive efficiency of the draft model.
Figure 1: In a single step of speculative decoding, the figure shows the probability of acceptance of draft tokens by their positions on GSM8K datasets, evaluated on FLAN-T5-small, base and large models. The draft model generates 30 tokens at each step.
Figure 2: Cascade Speculative Sampling uses LLMs of multiple sizes to generate and verify tokens. The model at the inner layer $\mathcal{M}_{d(i)}$ generates tokens that are proposed to and verified by LLMs at the next outer layer $\mathcal{M}_{d(i+1)}$. Once the verification process is completed, the Draft $\mathcal{M}_{d(i)}$ continues to generate new tokens given the accepted prefix tokens from previous layer $\mathcal{M}_{d(i-1)}$ as input.
Herein lies a trade-off: employing a larger draft model or generating more draft tokens per target run will improve the expected number of accepted tokens. However, either of these moves will increase the latency of the draft model significantly. As a result, the maximum speed improvement that the current speculative decoding algorithm can achieve is with an optimal balance between size and number of generated tokens. From previous work by Leviathan et al. (2023), it’s shown that very small models (less than 100 times the size of the target model) are optimal for most tasks experimented representing that the cost of drafting model can be the limiting factor of the speculative decoding algorithm. This observation leads to our central idea, due to the significant latency caused by the drafting model, improving drafting efficiency is the key to improving the performance of speculative decoding.
During our preliminary experiment, we found that speculative decoding with a simple non-neural language model can certainly improve the performance of a language model. This is because a simple non-neural language model is almost free of computational meaning that it’s free of the trade-off above. However, due to the limiting ability of such a simple non-neural language model, speculative decoding does not yield optimal results with a simple non-neural language model compared to a well-sized language model from the same model family of the target model. This makes a simple non-neural language model model suboptimal when a well-sized language model is available. We organize these observations into two laws of autoregressive generation of language models:
1. Without regard to generation length nor any additional hyperparameter tuning, a simple non-neural language model can always improve the generation speed of a language model by using speculative decoding over autoregressive generation.
2. When there is a well-sized smaller model(of the same architecture, distilled from the target model or trained from a similar corpus) and a task with enough expected generation length, the smaller model can be used to improve the generation speed of the target model over the simple non-neural language model.
These two laws provide us with insight into designing a new algorithm for even further speed improvement. We propose a new algorithm Cascade Speculative Drafting (CS. Drafting) a speculative execution algorithm with a paradigm that radically excludes any autoregressive generation from neural language models and exhaustively substitutes any simple non-neural language models with superior neural language models. The resulting algorithm contains multiple draft models with the smallest being a simple non-neural language model, with each neural draft model reviewing generations from a smaller model while proposing its reviewed content to a larger draft model or the target model. We name this layered speculative decoding design the Vertical Cascade.
Moreover, our experimental findings reveal an intriguing pattern: in a drafting step, the tokens generated later in the sequence by the draft model exhibit a progressively lower probability of being accepted by the target model, as illustrated in Figure 1. (Note that this is different from the acceptance rate which is the probability of the token being accepted conditioned on the previous token being accepted) The observation suggests that the later tokens generated by draft models, with a higher likelihood of rejection, contribute much less to the expected number of accepted tokens while creating the same latency. To address this, we propose using smaller, faster draft models $\mathcal{M}_{d(i-1)}$ for generating these high-rejection tokens, forming the Horizontal Cascade. Along with the aforementioned Vertical Cascade, these strategies compose our complete CS. Drafting approach illustrated in Figure 2. Integrating these vertical and horizontal cascades, our method innovatively mitigates LLM latency issues.
We evaluated our proposed CS. Drafting on a variety of tasks and model sizes. The experimental results demonstrate the superiority of our method. Moreover, theoretical analysis is provided to prove the optimality of our proposed method. Ablation studies further verify the effectiveness of our method.
In summary, our main contributions are as follows:
  • We introduce a novel algorithm, Cascade Speculative Drafting (CS. Drafting), a speculative-execution-based algorithm improving language models inference speed without compromising on generality or overall efficiency.
  • Our work includes a theoretical analysis affirming the optimality of our CS. Drafting approach.
  • We have established a new evaluation method, standardized walltime improvement, tailored for algorithms based on speculative decoding to ensure reproducibility and facilitate consistent comparisons across different studies.
  • Through our empirical study, we demonstrate that our CS. Drafting algorithm consistently achieves lower latency compared to vanilla speculative decoding across different tasks and settings. With a speedup of up to 72% over speculative decoding, this empirical evidence underscores the practical benefits and efficiency gains of our proposed approach.
2 Preliminary
2.1 Speculative Decoding Algorithm
In real-world applications of large language models (LLMs), the sequential nature of autoregressive decoding emerges as a primary latency bottleneck in terms of response time. Addressing this, speculative decoding (SD) (Leviathan et al., 2023) presents an innovative solution that utilizes a compact draft model for sequential token generation with validation by a larger target model resulting in reduced latency.
The core concept of speculative decoding involves a novel algorithm that accelerates sampling from autoregressive models without altering output distributions. At its heart, there are two key observations: 1) certain generations in language modeling are simpler than others and can be predicted by more efficient models correctly, and 2) using speculative execution along with a new sampling method enables faster, exact decoding from large models. generation
Specifically, let $x$ be the input tokens at a run and $\mathcal{M}_{t}$ and $\mathcal{M}_{d}$ are the target and the draft model respectively. We interpret speculative sampling as a two-stage operation. In the proposing stage, we sample $\{x_{1},...,x_{k}\}$ from draft model $\mathcal{M}_{d}$ autoregressively. In the reviewing stage, let $x_{i}\in\{x_{1},...,x_{k}\}$ represents the token at the current position, and we accept it if $\mathcal{M}_{d}(x)[i-1]<\mathcal{M}_{t}(x)[i-1]$ ; in the event that $\mathcal{M}_{d}(x)[i-1]>\mathcal{M}_{t}(x)[i-1]$, we reject $x_{i}$ with a probability of $1-\frac{\mathcal{M}_{t}(x)[i-1]}{\mathcal{M}_{d}(x)[i-1]}$ and proceed to resample $x_{i}$ from a recalibrated distribution $\text{norm}(\max(0,\mathcal{M}_{t}(x)[i-1]-\mathcal{M}_{d}(x)[i-1]))$ an then reject any token following $x_{i}$. Such a design guarantees the output is the same as sampling autoregressively using the target model alone.
This method was empirically validated on various tasks and model sizes, demonstrating a significant acceleration in inference times (2x-3x faster) compared to standard implementations, without affecting the outputs. Importantly, it does not require task-specific training, altering model architectures, or changing training procedures, making it a practical solution for enhancing the efficiency of autoregressive models such as Transformers (Leviathan et al., 2023).
2.2 Notation and Definition
We introduce notions and definitions that are used throughout the paper. Let $\mathcal{M}_{t}$ be the target model, $\{\mathcal{M}_{i}\}$ be the draft models assisting generation with $\mathcal{M}_{i}$ being the draft model generating the $i$-th token. In the simpler case of the speculative decoding, $\mathcal{M}_{i}=q$ for any $i$. Let $x$ be the input to the model at a single run, $\mathcal{M}_{t}(x)$ and $\mathcal{M}_{i}(x)$ are then the output probability distribution with input $x$.
Expected acceptance rate $\alpha=\mathbf{E}(1-\frac{1}{2}|\mathcal{M}_{t}(x)-\mathcal{M}_{i}(x)|)$ is the probability of draft generation by $\{\mathcal{M}_{i}\}$ being accepted by target model $\mathcal{M}_{t}$ (Leviathan et al., 2023).
Cost coefficient $c_{i}$ is the ratio of time for a single run of $\mathcal{M}_{i}$ over $\mathcal{M}_{t}$.
Lenience in speculative decoding is a parameter to loosen the review process by the target model for further speedup with the tradeoff that the generation result might not be the same as the target model. Yet in a recursive speculative decoding setup, lenience in the inner speculative decoding will not affect the final output. With lenience $l\in[1,\infty)$, The acceptance condition of $x_{i}$ speculative decoding will become $\mathcal{M}_{d}(x)[i]<l*\mathcal{M}_{t}(x)[i]$; otherwise, with a probability of $1-\frac{l*\mathcal{M}_{t}(x)}{\mathcal{M}_{d}(x)}$ , we reject the $x_{i}$ and any following tokens.
Expected walltime improvement factor (EWIF) is the expected time improvement of an algorithm based on speculative decoding over the vanilla generation of the target model.
3 Cascaded Speculative Drafting
Algorithm 1 CascadeSpeculativeDrafting
  Require:draft model $\{\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{n}}\}$, target model $\mathcal{M}_{t}$, $prefix$, flag $isFirstCall$, hyperparameters $K_{nn}$, $l$
  draftList $\leftarrow$ $[\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{n}}]$
  curGen $\leftarrow$ $prefix$, curProbs $\leftarrow$ []
   $\triangleright$ Unpack the current parameters from
  $[k_{1},...,k_{n-1}]\leftarrow$ first row of $K_{nn}$
   $\triangleright$ Base case when the MaG generate
  if draftList only contains one element then
     $\mathcal{M}\leftarrow$ first element of draftList
     $res\leftarrow$ $\mathcal{M}(curGen)$
      $\triangleright$ MaG generates greedily, so the logits will be 1 at the position of the generation token and 0 otherwise
     return $res.generation,res.logits$
  end if
   $\triangleright$ Begin the cascade
  for $i\leftarrow 1$ to $n$ do
     curTarget $\leftarrow$ the $i$-th item of draftList
     curDraftList $\leftarrow$ the sublist of draftList starting from index $i+1$
     curK $\leftarrow$ the submatrix of $K_{nn}$ from with left-top cornor (i + 1, j) to the right-most bottom
     while length of curGen is less than $k_{i}$ do
        $[x_{1},..,x_{r}],[p_{1},p_{2},...,p_{r}]\leftarrow$ CascadeSpeculativeDecoding(curDraftList, curTarget, curGen, curK, $l$)
        curGen $\leftarrow[x_{1},..,x_{s}]$
         $\triangleright$ Do not overwrite the previous probabilities
        s = curProbs.lengths + 1
        Add elements of $[p_{s},p_{2},...,p_{r}]$ to curProbs
     end while
  end for
   $\triangleright$ No leniency when the target model reviews
  if isFirstCall then
     $l=1$
  end if
   $\triangleright$ Review process by $\mathcal{M}_{t}$ to ensure the output distribution is the same as autoregressive generation
  $[p_{1}^{\prime},p_{2}^{\prime},...,p_{m}^{\prime}]\leftarrow$ $\mathcal{M}_{t}(curGen)$
  $index\leftarrow-1$
  for $i\leftarrow 1+prefix.length$ to $m-1$ do
     Sample $r$ from a uniform distribution
     if $r\leq min(1,\frac{lp_{i}^{\prime}}{p_{i}})$ then
        $index\leftarrow i$ and exit the loop
     end if
  end for
  if index is -1 then
     $index\leftarrow m$
  end if
  Sample $x_{last}$ from $p_{index}^{\prime}-p_{index}$
  return $prefix+[x_{1},...,x_{index-1}]+[x_{last}]$, $[p_{1}^{\prime},p_{2}^{\prime},...,p_{index}^{\prime}]$
3.1 Horizontal Cascade
One key observation that leads to our experiment is that during the drafting steps of speculative decoding, not all drafting tokens are made equal, as the first draft token is exponentially easier to be accepted than the $k$-th draft token since the expected acceptance rate of the $q_{i}$ decreases exponentially as $i$ increases. In Corollary 4.3, we proved the exponential decrease above and also observed that the derivative of EWIF with respect to $\alpha_{1}$ is also exponentially greater than the derivative of EWIF with respect to $\alpha_{k}$ representing the effort in improving model acceptance rate on the first token, though having the same cost, is exponentially more valuable than the effort on the last token. As a result, given multiple models available for drafting, we propose cascade speculative decoding which involves using a larger model to generate the first few tokens and a smaller model to generate the remaining tokens.
In addition, assuming $\mathcal{M}_{d1},\mathcal{M}_{d2}$ being two draft models [^1] The notation of $\mathcal{M}_{d1}$ is different from $\mathcal{M}_{i}$ When the $d$ appeared in the subscript, it means the model can be used for generating multiple tokens. with $\mathcal{M}_{d1}$ having high cost and acceptance rate then $\mathcal{M}_{d2}$ during drafting of $i$-th token with cascade decoding, a more general setup can be achieved by using an interpolated draft model by assigning a probability $f(i)$ such that with probability $f(i)$ the $i$-th token is generated by $\mathcal{M}_{d1}$; otherwise, it’s generated by $\mathcal{M}_{d2}$. This interpolated draft model setup can provide a smoother transition of cost and acceptance rate. However, we prove that such a setup will not improve the walltime improvement factor in theorem 4.4.
3.2 Vertical Cascade
With our proven results of theorem 4.4 that for optimal cascade sampling of two draft model $\mathcal{M}_{d1},\mathcal{M}_{d2}$, each draft token will be generated by one of the draft models, we observe one additional inefficiency even with the already improved efficiency with horizontal cascade. As we have two draft models $\mathcal{M}_{d1},\mathcal{M}_{d2}$ with $\mathcal{M}_{d1}$ having better acceptance rate and higher time cost, we hypothesize that another round of speculative sampling replacing autoregressive generation of $\mathcal{M}_{d1}$ for drafting can further improve the efficiency of cascade sampling. Furthermore, leniency introduced by Leviathan et al. (Leviathan et al., 2023) provides a tradeoff between quality and speed. In practice, such a quality drop can be undesirable; however, the Cascade Speculative Drafting algorithm does not have such a limitation. Since the final quality is always guaranteed by the target model, any lenience in speculative decoding between the draft models will not affect the final output. In addition, we found that lenience can potentially bring further speedup as the optimal matching to the larger draft model in speculative drafting might not reduce the generation speed as much as the speedup introduced by lenience.
3.3 Max-gram for Better simple non-neural language model
A simple non-neural language model is special to speculative decoding as it almost incurs no cost when drafting for its target model. This means despite a slight implementation overhead, it can almost always replace autoregressive generation using greedy decoding, sampling, or beam search. In addition, in the CS Drafting algorithm, a simple non-neural language model also generates for the draft model that reviews the most tokens making the output alignment between the simple non-neural language model and the smallest draft model very important for the speedup of our algorithm.
During our pursuing of a better simple non-neural language model, we noticed a general pattern. When language models generate, some words and phrases in the input query appear repeatedly in the generation result. A common sense that the answer should have words and phrases overlapping with the question.
Based on the observation, we designed the Max-Gram(MaG) algorithm that greedily finds maximal matches in the initial input(or existing generation) and tokens from the end of the generation, and in case there is no match, we use a bigram model pretrained on the Wikipedia. (We chose Wikipedia to not affect the generality of our experiment) To our surprise, despite its simplicity, we found this simple and general algorithm very effective in aligning with language model model output. We included a GPU-friendly version of the Max-Gram algorithm in the appendix A.
3.4 Algorithm
Combining the horizontal and vertical cascades, the algorithm is shown in Algorithm 1. The center of it has the a for loop which performs the horizontal cascade while the recursive call forms the vertical cascade. Noticeably, the algorithm with the last element of the draft models being the MaG algorithm, there will be no neural language models generate autoregressively.
4 Theoretical Analysis
We perform theoretical analysis to provide simple heuristics for evaluating a CS. Drafting system, and to support the generality of CS. Drafting, we prove the optimality of horizontal cascade.
4.1 Expected Walltime Improvement
We provide the formula for EWIF in an extended setting when multiple draft models are involved.
Theorem 4.1.
the expected walltime improvement factor(EWIF) by Algorithm 1 is $T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\sum_{i=0}^{k}\prod_{j=1}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$
Proof.
Since every token will only be considered by target model if all tokens before are accepted, a simple recursion leads to the probability of acceptance of the $i$-th token being $\prod_{j=1}^{i-1}\alpha_{j}$, and the $\prod_{j=1}^{0}\alpha_{j}=1$ representing the token generated by the target model. In the formula of the EWIF, the numerator is the expected number of tokens generated. The denominator is the time costs for each step of such a system. The formula is then the EWIF. ∎
We can immediately achieve a formula for speculative decoding by requiring all the draft models to be the same.
Corollary 4.2.
When $q_{1}=q_{2}=...=q_{k}$, let $\alpha_{i}=\alpha$ and $c_{i}=c$ for any ${}_{1}$, the EWIF is $\frac{1-\alpha^{k+1}}{(1-\alpha)(ck+1)}$
Furthermore, theorem 4.1 can be used to analyze the importance of the tokens in the drafting step.
Corollary 4.3.
The probability of $i$-th token being accepted is $\prod_{j=1}^{i}\alpha_{j}$. The derivative of EWIF with respect to $\alpha_{l}$ is $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{l}}=\frac{\sum_{i=l}^{k}\prod_{j=1,j\neq l}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$. Specifically, $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{1}}=\frac{\sum_{i=1}^{k}\prod_{j=2}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$ and $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{k}}=\frac{\prod_{j=1}^{k-1}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$
Using the information provided by (Leviathan et al., 2023), we calculated a simulated EWIF with the assumption that the event of acceptance by the target model is a Bernoulli trial. Results are shown in table 1 which shows that the cascade speculative sampling achieved better EWIF over vanilla speculative under the assumption.
4.2 Generality of Horizontal Cascade
The setting of cascade drafting can be simply extended using model interpolation that involves choosing between two models using randomness. Given two draft models $d_{1}$ and $d_{2}$ and a probability function $f$, $q_{i}$ is considered to be $f{i}d_{1}+(1-f(i))d_{2}$i.e with probability $f(i)$, $q_{i}$ is $d_{1}$, with probability $1-f{i}$ $q_{i}$ is $d_{2}$, as a result, $\alpha_{i}=f(i)\alpha_{d1}+(1-f(i))\alpha_{d2}=\alpha_{d2}+f(i)(\alpha_{d1}-\alpha_{d2})$ and $c_{i}=f(i)c_{d1}+(1-f(i))c_{d2}=c_{2}+f(i)(c_{d1}-c_{d2})$. In theorem 4.4, we prove that the optimal can always be reached without interpolation.
Theorem 4.4.
Given an cascade drafting system with the setting mentioned above. There exists a probability function $f$ that optimizes the system such that $f$ satisfies $f(i)\in\{0,1\}$ for any $i$. This means each of the tokens will be generated by either $d_{1}$ or $d_{2}$, and the optimal system is deterministic.
Proof of Theorem 4.4 is included in appendix B
Dataset $\mathcal{M}_{d1}$ $\mathcal{M}_{d2}$ $k$ $k_{2}$ EWIF
CNNDM T5-small - 9 - 2.65
CNNDM T5-base - 8 - 2.96
CNNDM T5-base T5-small 5 3 3.03
ENDE T5-small - 12 - 3.61
ENDE T5-base - 11 - 3.75
ENDE T5-base T5-small 5 8 3.93
Table 1: Simulated EWIF under the assumption that the acceptance distribution is a Bernoulli distribution. In the simulation, cascade drafting exceeded the performance of the speculative decoding on both CNN Dailymail and WMT EnDe datasets.
5 Evaluation
Previous works on speculative decoding and related methods relied on walltime as an evaluation method. However, there are standardization and legitimacy concerns related to walltime. Leviathan et al. (2023) mentioned that cost coefficiency, the ratio between the time for a single run of draft model and target model will vary from system to system which can result in significant variation in wall time improvement across different systems, making comparisons across different research work difficult.
In addition, a recent analysis suggests GPU speed can vary for the same GPU model with one being 1.5x faster than another GPU of the same model (Sinha et al., 2022), and even for the same GPU, the performance variation between different runs is unknown; therefore, theoretical walltime improvement up to 1.5x can be undermined by GPU variation, and methods without any improvement can be measured with an improvement up to 1.5x by running on a different GPU. The properties of such lack of standardization and stability make walltime unfit for benchmarking algorithmic improvement. To resolve this issue, we designed a new benchmarking method to measure speculative decoding and its related algorithms which can reduce the effect of different/unstable GPU performance.
Our proposed method, standardized walltime improvement(SWI) calculates the GPU times of the models, assuming that each run of a language model costs the same amount of time, an assumption made when inventing the speculative sampling algorithm (Chen et al., 2023). Given a speculative decoding system $S$ containing k draft models $\mathcal{M}_{d1},...,\mathcal{M}_{dk}$, target model $\mathcal{M}_{t}$ solving a task $T$, and let $count(S,m,T)$ be the count of runs of model $m$ during the execution of task $T$ with system $S$, let $c_{i}$ be the expected cost coefficient being the time ratio between a single run of $\mathcal{M}_{i}$ and $\mathcal{M}_{t}$, and let $n$ be the total of tokens generated, the standardized walltime improvement is defined as
$\begin{split}SWI&(\mathcal{M}_{d1},...,\mathcal{M}_{dk},\mathcal{M}_{t},T)=\\ &\frac{n}{count(S,\mathcal{M}_{t},T)+\sum_{i=1}^{n}count(S,\mathcal{M}_{di},T)c_{i}}\end{split}$
which is simply the sum of GPU time by all models when $c_{i}$ is measured during a local experiment. The benefit of the $SWI$ evaluation is that $count(S,\mathcal{M}_{t},T)$ and $count(S,\mathcal{M}_{di},T)$ will be the same across different systems, so by using the same $c_{i}$ across different experiments the GPU performance variation will be eliminated; therefore, it provides a standardized, reproducible benchmark.
During our experiment, we will use two sets of $\{c_{i}\}$, existing $\{c_{i}\}$ reported by (Leviathan et al., 2023) and an heuristic $\{c_{i}\}$ being the ratio of number of tunable parameters between $\mathcal{M}_{di}$ and $\mathcal{M}_{t}$.
Dataset Algorithm $\{\mathcal{M}_{di}\}$ SWI (MS) $k_{11}$ $k_{21}$ $k_{12}$ $l$
GSM8K Autoregressive - 1 - - - -
GSM8K S. Decoding base $3.38$ 10 - - -
GSM8K S. Decoding small $3.06$ 11 - - -
GSM8K C.S. Drafting base, mag $3.62$ 9 - - -
GSM8K C.S. Drafting small, mag $3.16$ 11 - - -
GSM8K C.S. Drafting base, small, mag 3.75 7 1 3 1.5
MMLU Autoregressive - 1 - - - -
MMLU S. Decoding base $3.97$ 13 - - -
MMLU S. Decoding small $4.12$ 19 - - -
MMLU C.S. Drafting base, mag $4.50$ 12 - - -
MMLU C.S. Drafting small, mag $4.36$ 17 - - -
MMLU C.S. Drafting base, small, mag 4.83 5 1 14 2
Dataset Algorithm $\{\mathcal{M}_{di}\}$ SWI (PW) $k_{11}$ $k_{21}$ $k_{12}$ $l$
GSM8K Autoregressive - 1 - - - -
GSM8K S. Decoding base $2.99$ 8 - - -
GSM8K S. Decoding small $2.76$ 8 - - -
GSM8K C.S. Drafting base, mag $3.22$ 8 - - -
GSM8K C.S. Drafting small, mag $2.90$ 7 - - -
GSM8K C.S. Drafting base, small, mag 3.29 5 1 4 1.5
MMLU Autoregressive - 1 - - - -
MMLU S. Decoding base $3.42$ 10 - - -
MMLU S. Decoding small $3.51$ 11 - - -
MMLU C.S. Drafting base, mag $4.07$ 9 - - -
MMLU C.S. Drafting small, mag $3.97$ 12 - - -
MMLU C.S. Drafting base, small, mag 4.23 5 1 8 1.5
Table 2: The experiment results from dataset GSM8K and MMLU. BASE and SMALL refer to FLAN-T5-base and FLAN-T5-small. SWI(model size) is the standardized walltime improvement with the assumption the latency of each run of a model is its number of parameters(MS) and time cost data from previous work(PW)(Leviathan et al., 2023) respectively.
6 Experiment
6.1 Setup
We choose our target and draft models from FLAN-T5 (Chung et al., 2022) family for our experiment as there are large variations in model sizes in the FLAN-T5 family (from 77 million to 11 billion); also, the reasoning ability of the FLAN-T5 models providing more options for our experiments. To make our experiment similar to real-world usage, We do not perform fine-tuning and the generation is done in a zero-shot manner. We use FLAN-T5-xxl as our target model, and T5-base and T5-small as our reviewing draft models and Max-Gram algorithm to be the generating draft model.
With limited computational budge, we are only able to perform sampling in temperate 00 (greedy decoding) due to the extra steps of hyperparameter tuning needed for the Cascade Speculative Decoding algorithm; the reason behind this is the benefit of temperate 00 that we can reuse the target model output which is deterministic which saves a great amount of computational resource on running the target models repeatedly.
Cascade Speculative Drafting Inference To simplify the number of hyperparameters to tune, we use MaG to generate 10 tokens at once as it’s rare that more than 10 tokens will be accepted, and we do not use lenience when the reviewer is $\mathcal{M}_{t}$ to ensure the output distribution remains do not change, and we also eliminate lenience between $\mathcal{M}_{d2}$ and MaG due to the performance map between them are large.
With these constraints, we are left with at most four parameters $k_{11}$, $k_{21}$, $k_{12}$, $l$. For the step of CS Drafting when the target model is the reviewer $k_{11}$ and $k_{12}$ are used. $k_{21}$, $l$ in the step when $\mathcal{M}_{d1}$ is the reviewer. However, there are still too many hyperparameters for us to search for the optimal result exhaustively; instead, we use halving grid search to search for a set of hyperparameters to compare to the baseline.
Baseline We consider speculative decoding and vanilla autoregressive generation as our baseline. For speculative decoding, we consider T5-small and T5-base as the draft model as they perform better than T5-large in previous work(Leviathan et al., 2023) and our preliminary experiment. We try a large range of k-values from 2 to 30 to ensure that we will find the best result for each model as the draft model to serve as our baseline.
Dataset We chose reasoning tasks for our experiment as they present more challenges of real-world NLP applications than summarization tasks. For both datasets, we experiment in a zero-shot chain-of-thought (Wei et al., 2023) setup.
The datasets we experimented on are the following:
GSM8k (Cobbe et al., 2021) A challenging mathematical reasoning dataset many large language models struggle with.
MMLU (Hendrycks et al., 2021) A reasoning dataset in many different subjects. MMLU does not have training sets meaning applying task-specific methods such as knowledge distillation to improve speculative decoding can be challenging.
6.2 Experimental Results
The experiments are shown in table 2. In two settings of SWI and Walltime, Cascade Speculative Decoding has exceeded the speculative decoding algorithm. For GSM8K, Cascade Speculative Decoding has a maximum improvement of $37\%$ over the fastest speculative algorithm; for MMLU, the maximum improvement of Cascade Speculative Decoding over speculative decoding is $72\%$.
One other observation is that despite during speculative decoding t5-small has out-performed t5-base on some experiments, but in CS Drafting with only one neural model and MaG, T5-base consistently outperforms t5-small.
7 Related Work
7.1 Efficienct Methods for Natural Language Inference
In the era of large language models, the efficiency during inference becomes a key to model service. To reduce the model inference cost and speed up, several efficient methods have been proposed, including pruning, knowledge distillation and quantization (Treviso et al., 2023). Model pruning takes structured (Xia et al., 2022; Voita et al., 2019) or unstructured (Guo et al., 2021; Chen et al., 2020) methods to remove the redundant model parameters to reduce the storage memory and increase inference speed. Knowledge distillation takes the approach of transferring knowledge from a superior teacher model to a smaller student model (Hinton et al., 2015; Gou et al., 2021). Quantization maps high-precision data representations (e.g. 32 bits) into low-precision ones (e.g. 8 bits) to reduce memory consumption (Bhandare et al., 2019; Schaefer et al., 2023).
7.2 Speculative Decoding
Recent works improved the speed of speculative decoding by experimenting with task-specific knowledge distillation (Zhou et al., 2023), or using approaches similar to speculative decoding to achieve a tradeoff between quality and speed (Leviathan et al., 2023; Kim et al., 2023). However, the reduced quality for speed can be undesirable, and task-specific training might not be applicable to many real-world NLP applications such as chatbots; overall, we found no generic improvement for speculative decoding in the current literature.
Speculative decoding (Chen et al., 2023; Leviathan et al., 2023) is proposed to improve inference speed by using a draft model to perform auto-regressive decoding and a target model to check and replace the generated tokens. A key to further speed up is to increase the acceptance rate in speculative decoding, which can be done by aligning the distribution between the draft model and the target model. Some recent work has proposed to improve the performance of speculative decoding in a domain/task-specific setting. Zhou et al. (2023) use generalized knowledge distillation to finetune the draft model and achieve 10%-45% speedup compared to draft models without knowledge distillation. Liu et al. (2023) propose Online Speculative Decoding to continually update multiple draft models on observed user query data, and the draft models align with the query distribution in real-time. However, both approaches require additional task-specific training making it difficult to apply the methods in a general setting. On the other hand, He et al. (2023) proposed constructing a datastore from a domain-specific corpus to serve as a draft model which can be used when a draft model is not available.
8 Conclusion
In our work, we proposed a novel algorithm CS. Drafting that involves two cascades, the vertical cascade, and the horizontal cascade. The vertical cascade eliminates the necessity of autoregressive generation from a neural language model, and the horizontal cascade effectively allocates the cost for drafting tokens at different positions. Our theories proved the optimality of the horizontal cascade, and empirically, we showed that CS. Drafting achieved up to 72 percent more speedup over speculative decoding without any additional training while maintaining generation quality.
Appendix A Max-Gram Implementation
Listing 1: Max-Gram Algorithm
1def torch_index(t, value):
2    return (t == value).nonzero(as_tuple=True)[0][0]
3
4def max_gram(input_ids, encoder_ids, n=1):
5    matches = (encoder_ids[0] == input_ids[0, -1]).int()
6    if matches.sum() < 1:
7        return None
8    for i in range(2, input_ids.shape[-1] + 1):
9        new_matches = (encoder_ids[0, :(-1 * (i - 1))] == input_ids[0, -1 * i]).int()
10        combined_matches = (2 - new_matches == matches[1:]).int()
11        if combined_matches.sum() < 1:
12            index = torch_index(torch.cat(
13                (
14                    torch.tensor([0] * (i - 1), device=torch.device(encoder_ids.device)),
15                    matches
16                ),
17                dim=-1
18            ), 1)
19            return encoder_ids[:, index:index + n]
20        else:
21            matches = combined_matches
22    index = torch_index(torch.cat((
23        torch.tensor([0] * (encoder_ids.shape[-1] - matches.shape[-1])), matches), dim=-1
24        ), 1)
25    return encoder_ids[:, index+1:index + n+1]
Appendix B Proof and Analysis of Theorem 4.4
We first prove a lemma that helps the proof of theorem 4.4
Lemma B.1.
$\forall a,b,c,d\in\mathbb{R}_{>0},\gamma\in(0,1)$
$\frac{\gamma a+(1-\gamma)b}{\gamma c+(1-\gamma)d}\leq\max(\frac{a}{c},\frac{b}{d})$
Proof.
Without loss of generality, assume $\frac{a}{c}\geq\frac{b}{d}$,
$\frac{a}{c}\geq\frac{\gamma a+(1-\gamma)b}{\gamma c+(1-\gamma)d}$
$\Longleftrightarrow a(\gamma c+(1-\gamma)d)\geq c(\gamma a+(1-\gamma)b)$
$\Longleftrightarrow(1-\gamma)ad\geq(1-\gamma)cb$
$\Longleftrightarrow\frac{a}{c}\geq\frac{b}{d}$
Proof of Theorem 4.4
Proof.
Let $\alpha_{d_{1}},\alpha_{d2}$ and $c_{d_{1}},c_{d2}$ be the acceptence rate and cost ratio for model $d_{1},d_{2}$ respectively. Without loss of generality, we assume $\alpha_{d_{1}}>\alpha_{d2}$ and $c_{d_{1}}>c_{d2}$ Then the system satisfies $\alpha_{i}=f(i)\alpha_{d_{1}}+(1-f(i))\alpha_{d_{2}}$ and $c_{i}=f(i)\alpha_{d_{1}}+(1-f(i))c_{d_{2}}$ We prove using contradiction. Assuming all optimal systems need at least one interpolation of models at some positions. Let an optimal system of $f$ given $d_{1},d_{2}$, Without loss of generality, assume $r$ to be the first indices such that $f(r)$ not being an integer and the system $f$ to be the one with largest $r$ among all optimal systems.
$T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\alpha_{r}(1+\sum_{i=r+1}^{k}\prod_{j=r+1}^{i}\alpha_{j})+\sum_{i=0}^{r-1}\prod_{j=1}^{i}\alpha_{j}}{1+c_{r}+\sum_{i=1,i\neq r}^{k}c_{i}}$
$=\frac{(f(r)\alpha_{d_{1}}+(1-f(r))\alpha_{d_{2}})(1+\sum_{i=r+1}^{k}\prod_{j=r+1}^{i}\alpha_{j})+\sum_{i=0}^{r-1}\prod_{j=1}^{i}\alpha_{j}}{1+f(r)c_{d_{1}}+(1-f(r))c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i}}$
Let $A_{1}=1+\sum_{i=r+1}^{k}\prod_{j=r+1}^{i}\alpha_{j},A_{2}=\sum_{i=0}^{r-1}\prod_{j=1}^{i}\alpha_{j}$
$=\frac{f(r)(\alpha_{d_{1}}A_{1}+A_{2})+(1-f(r))(\alpha_{d_{2}}A_{1}+A_{2})}{f(r)(1+c_{d_{1}}+\sum_{i=1,i\neq r}^{k}c_{i})+(1-f(r))(1+c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i})}$
$\leq\max{\frac{(\alpha_{d_{1}}A_{1}+A_{2})}{1+c_{d_{1}}+\sum_{i=1,i\neq r}^{k}c_{i}},\frac{\alpha_{d_{2}}A_{1}+A_{2})}{1+c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i})}}$
Meaning that $f(r)=0$ or $f(r)=1$ will result in a time improvement factor no less than a non-integer $f$, meaning replacing $f(r)$ with an integer will still result in an optimal system. This contradicts with the assumption that $f$ is the system with the largest $r$. As a result, there exists an optimal system without interpolation. ∎
Figure 3: Acceptance rate of draft tokens by their positions on MMLU datasets, evaluated on FLAN-T5-small, base and large models.