### v2

# Cascade Speculative Drafting for Even Faster LLM Inference

Abstract
Speculative decoding reduces the latency of large language model (LLM) inference by leveraging a draft model to draft for a larger target model to review. However, drafting in speculative decoding involves slow autoregressive generation and allocates the same amount of time to generating tokens of varying importance. These two inefficiencies lead to its suboptimal performance. To address this issue, we introduce Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises two types of cascades. The Vertical Cascade eliminates autoregressive generation from neural models, while the Horizontal Cascade constitutes efficient time allocation in drafting. Combining both cascades, our CS Drafting algorithm has achieved up to a 72 percent additional speedup over speculative decoding in our experiments, while maintaining the same output distribution as the target model.
Machine Learning, ICML
[…]
1 Introduction
The advent of Large Language Models (LLMs), like GPT-4 (OpenAI, 2023) has marked a significant milestone in the field of natural language processing (NLP). These models have not only excelled in various NLP tasks but have also found widespread application in user-interactive settings such as chatbots and virtual assistants. This application involves an extremely high number of users, up to hundreds of millions daily. To serve in real-time at this scale, a low-latency system is not only cost-saving but also crucial for keeping the service running. In addition, the sheer scale of the service means that even a slight improvement in the latency of LLMs can greatly contribute to both the provider and the community. Consequently, optimizing LLM latency has become a crucial and impactful research focus.
Unfortunately, LLMs currently suffer from high latency, especially in long-form generation, as autoregressive LLMs generate tokens one by one. An emerging solution, known as speculative decoding (Leviathan et al., 2023; Chen et al., 2023), shows potential to mitigate this issue. In speculative decoding, a draft model (which is smaller and faster) generates $k$ tokens in each step (with $k$ being a hyperparameter), and these tokens are then reviewed by a target model (which is larger and slower) in parallel. The target model will accept any tokens aligned with its output and further generate one token; the process happens all in a single run while maintaining the output distribution identical to the autoregressive one by the target model itself. As a larger number of expected accepted tokens per run indicates fewer target model runs, the efficiency of speculative decoding can be attributed to two main factors: the expected number of tokens accepted by the target model per run, and the latency of drafting.
Figure 1: The probability of acceptance of draft tokens in relation to their positions in a single step of speculative decoding, evaluated on FLAN-T5-small, base, and large models. The draft model generates 30 tokens at each step.
Figure 2: Cascade Speculative Sampling uses LLMs of multiple sizes to generate and verify tokens. The model at the inner layer $\mathcal{M}_{d(i)}$ generates tokens proposed to and verified by LLMs at the next outer layer, $\mathcal{M}_{d(i+1)}$. This recursive process forms the vertical cascade. After a vertical cascade, a smaller $\mathcal{M}_{d(i)}$ will serve as the target model and perform a vertical cascade. The process repeats until the smaller model becomes a statistical language model that will generate autoregressively forming the horizontal cascade.
[…]
In light of this, there exist a type of draft models, including statistical language models like bigram language models, that are immune to drafting efficiency issues. Owing to their simple structure, these models incur negligible latency and computational resource costs compared to neural language models, thereby avoiding the aforementioned trade-off. Therefore, assuming enough parallel computing power to perform computation on additional $k$ tokens with similar time cost (Leviathan et al., 2023), utilizing speculative decoding with a statistical language model as the draft model can generally and intuitively enhance the efficiency of a neural language model. However, due to their limited capabilities, speculative decoding with statistical language models alone may not yield optimal results compared to using a well-sized neural language model from the same family as the target model when the neural language model is available.
Based on the above analysis, we propose Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises multiple draft models, starting with the smallest being a statistical language model. Each neural draft model reviews generations from a smaller model and then proposes its reviewed content to either a larger draft model or the target model. We refer to this tiered speculative decoding approach as the Vertical Cascade.
Furthermore, our analysis in Figure 1 reveals an interesting pattern in the expected acceptance rate during the drafting step: tokens generated later in the sequence by the draft model show a progressively lower probability of being accepted by the target model [^1] A theoretical analysis is provided in Section 4. Note that this probability differs from the acceptance rate, which is the probability of a token being accepted conditioned on the previous tokens’ acceptance.. This indicates that later tokens from draft models, more prone to rejection, contribute less to the expected number of accepted tokens, yet incur the same latency. Based on this observation and our practice, we suggest the use of smaller, faster draft models for generating these high-rejection tokens, forming the Horizontal Cascade. Along with the aforementioned Vertical Cascade, these strategies compose our complete CS Drafting approach, as illustrated in Figure 2.
Through theoretical analysis and empirical studies, we demonstrate that the CS Drafting algorithm outperforms speculative decoding in terms of latency across various tasks and settings, achieving an additional speedup of up to 72% over speculative decoding. These findings highlight the practical advantages and efficiency enhancements offered by both vertical and horizontal cascades.
Our main contributions are summarized as follows:
  • We introduce Cascade Speculative Drafting (CS Drafting), a speculative-execution-based algorithm that improves language model inference speed without sacrificing generation quality.
  • We provide theoretical analyses supporting the effectiveness of the proposed CS Drafting approach.
  • We conduct empirical experiments showing that CS Drafting achieves further speedup over speculative decoding across different tasks and settings.
2 Preliminary
2.1 Speculative Decoding Algorithm
In real-world applications of large language models (LLMs), the sequential nature of autoregressive decoding emerges as a primary latency bottleneck in terms of response time. Addressing this, speculative decoding (SD) (Leviathan et al., 2023) presents an innovative solution that utilizes a compact draft model for sequential token generation with validation by a larger target model resulting in reduced latency.
The core concept of speculative decoding involves a novel algorithm that accelerates sampling from autoregressive models without altering output distributions. At its heart, there are two key observations: 1) certain generations in language modeling are simpler than others and can be predicted by more efficient models correctly, and 2) using speculative execution along with a new sampling method enables faster, exact decoding from large models.
Specifically, let $x$ be the input tokens at a run and $\mathcal{M}_{t}$ and $\mathcal{M}_{d}$ are the target and the draft model respectively, and $\mathcal{M}_{t}(x)[i]$ and $\mathcal{M}_{d}(x)[i]$ be their probability output at $i$-th token when input is $x$. We interpret speculative sampling as a two-stage operation. In the proposing stage, we sample $\{x_{t+1},...,x_{t+k}\}$ from draft model $\mathcal{M}_{d}$ autoregressively and append them to $x$ In the reviewing stage, let $x_{i}\in\{x_{t+1},...,x_{t+k}\}$ represents the token at the current position, and we accept it if $\mathcal{M}_{d}(x)[i-1]<\mathcal{M}_{t}(x)[i-1]$; in the event that $\mathcal{M}_{d}(x)[i-1]>\mathcal{M}_{t}(x)[i-1]$, we reject $x_{i}$ with a probability of $1-\frac{\mathcal{M}_{t}(x)[i-1]}{\mathcal{M}_{d}(x)[i-1]}$ and proceed to resample $x_{i}$ from a recalibrated distribution $\text{norm}(\max(0,\mathcal{M}_{t}(x)[i-1]-\mathcal{M}_{d}(x)[i-1]))$ and reject any token following $x_{i}$. At the end, the target model will generate one additional token following the accepted tokens. Such a design guarantees the output is the same as sampling autoregressively using the target model alone (Leviathan et al., 2023).
This method was empirically validated on various tasks and model sizes, demonstrating a significant acceleration in inference times (2x-3x faster) compared to standard implementations, without affecting the outputs. Importantly, it does not require task-specific training, altering model architectures, or changing training procedures, making it a practical solution for enhancing the efficiency of LLMs.
3 Cascaded Speculative Drafting
Algorithm 1 CascadeSpeculativeDrafting
  Require:draft model $\{\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{n}}\}$, target model $\mathcal{M}_{t}$, $prefix$, flag $isFirstCall$, hyperparameters $K_{nn}$, $l$
  draftList $\leftarrow$ $[\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{n}}]$
   $\triangleright$ Initialize curGen and curProb
  curGen $\leftarrow$ $prefix$, curProbs $\leftarrow$ a list of ones of length prefix
   $\triangleright$ Unpack the a list of $k$ for current function call
  $[k_{1},...,k_{n-1}]\leftarrow$ first row of $K_{nn}$
   $\triangleright$ Base case of the recursive call when the MaG generate
  if draftList only contains one element then
     $\mathcal{M}\leftarrow$ first element of draftList
     $res\leftarrow$ $\mathcal{M}(curGen)$
      $\triangleright$ MaG generates greedily, so the logits will be 1 at the position of the generation token and 0 otherwise
     return $res.generation,res.logits$
  end if
   $\triangleright$ Begin the cascade
  for $i\leftarrow 1$ to $n$ do
     curTarget $\leftarrow$ the $i$-th item of draftList
     curDraftList $\leftarrow$ the sublist of draftList starting from index $i+1$
     curK $\leftarrow$ the submatrix of $K_{nn}$ from with the top-left corner at $(i+1,j)$ extending to the bottom-right corner
     while length of curGen - prefix.length is less than $k_{i}$ do
        $[x_{1},..,x_{r}],[p_{1},p_{2},...,p_{r}]\leftarrow$ CascadeSpeculativeDecoding(curDraftList, curTarget, curGen, False, curK, $l$)
        curGen $\leftarrow[x_{1},..,x_{r}]$
         $\triangleright$ Do not overwrite the previous probabilities
[…]
        Add elements of $[p_{s},p_{2},...,p_{r}]$ to curProbs
     end while
  end for
   $\triangleright$ No leniency when the target model reviews
  if isFirstCall then
     $l\leftarrow 1$
  end if
   $\triangleright$ Review process by $\mathcal{M}_{t}$ to ensure the output distribution is the same as autoregressive generation
  $[p_{1}^{\prime},p_{2}^{\prime},...,p_{m}^{\prime}]\leftarrow$ $\mathcal{M}_{t}(curGen)$
  $index\leftarrow m$
  for $i\leftarrow 1+prefix.length$ to $m-1$ do
     Sample $r$ from a uniform distribution
     if $r\leq min(1,\frac{lp_{i}^{\prime}}{p_{i}})$ then
        $index\leftarrow i$ and exit the loop
     end if
  end for
  Sample $x_{last}$ from $p_{index}^{\prime}-p_{index}$
  return $[x_{1},...,x_{index-1}]+[x_{last}]$, $[p_{1}^{\prime},p_{2}^{\prime},...,p_{index}^{\prime}]$
3.1 Vertical Cascade
The speculative decoding algorithm relies on the autoregressive generation of a smaller draft model. However, despite its smaller size, the draft model must run $k$ times for each target model run, and the cost can still be significant. To further reduce latency in the drafting process, we propose using an even smaller model to assist in drafting and employing the original draft model to review the generation of this smaller model. This process can be done recursively until it reaches the smallest draft model. The smallest draft model can be a statistical language model, such as a bigram language model, where the cost is negligible compared to neural language models. We refer to this recursive process as Vertical Cascade.
Furthermore, to further improve speed, we can introduce leniency, a hyperparameter in speculative decoding that loosens the review process by the target model, allowing for faster speed at the trade-off of potentially differing results from the target model (Leviathan et al., 2023). With leniency $l\in[1,\infty)$, the acceptance condition for token $x_{i}$ in speculative decoding becomes $\mathcal{M}{d}(x)[i]<l\times\mathcal{M}{t}(x)[i]$. Otherwise, with a probability of $1-\frac{l\times\mathcal{M}{t}(x)}{\mathcal{M}{d}(x)}$, we reject $x_{i}$ and any following tokens. While the reduced quality introduced by leniency in speculative decoding is generally undesirable in practice, it fortunately does not affect the final output in our vertical cascade approach—we can only apply leniency in the inner speculative decoding. As long as leniency is not applied to the target model, the final quality is always guaranteed by the target model.
3.2 Horizontal Cascade
As illustrated in Figure 1, another key observation is that during the drafting steps of speculative decoding, not all drafting tokens are created equal. The first draft token is more likely to be accepted as it only depends on itself; the last token is rarely accepted, as it has a chance of being reviewed only if all preceding tokens are accepted. From another perspective, one could also say that the first draft token bears the responsibility for all succeeding tokens, while the last token bears none. In Section 4, we provide a theoretical analysis of this phenomenon.
Based on this observation, we propose better allocation of the time cost of drafting by using larger models to generate the preceding draft tokens and smaller models for the succeeding tokens. This design allows more time to be allocated to more effective tokens, leading to a reduction in overall latency. We refer to this design as the Horizontal Cascade.
3.3 Max-gram for Better Statistical Drafting
A statistical language model is special to speculative decoding, as it incurs almost no cost when drafting for its target model. In our pursuit of a more effective statistical language model, we noticed a general pattern: in language model generation, some words and phrases from the input query frequently reappear in the generated content. Based on this observation, we designed the Max-Gram (MaG) algorithm. It greedily identifies maximal matches between the initial input (or existing generation) and tokens from the end of the generation. In cases where there is no match, we resort to a bigram model based on the probability distribution of Wikipedia (chosen to maintain the generality of our experiment). Surprisingly, despite its simplicity, this straightforward and general algorithm has proven highly effective in aligning with neural language model outputs. We include a GPU-friendly version of the Max-Gram algorithm in Appendix A.
3.4 Algorithm
Combining the horizontal and vertical cascades, the algorithm is presented in Algorithm 1. At its center, there is a for-loop that performs the horizontal cascade, while the recursive call forms the vertical cascade. Notably, the last element of the draft models is the MaG model.
The algorithm requires a two-dimensional hyperparameter, $K_{nn}$, with each row serving as the stop criteria for a layer of recursive calls. For simplicity, we assume the leniency parameter $l$ is universal for the algorithm, except when the target model is under review; thus, the algorithm can benefit from the speedup of leniency without altering the output distribution.
4 Theoretical Analysis
In this section, we present some theoretical analyses of the proposed Cascade Speculative Drafting algorithm.
4.1 Expected Walltime Improvement
To assist our theoretical analysts, we establish the notions. Let $\mathcal{M}_{t}$ be the target model, $\{\mathcal{M}_{i}\}$ be the draft models assisting generation with $\mathcal{M}_{i}$ being the draft model generating the $i$-th token. In the simpler case of the speculative decoding, $\mathcal{M}_{i}=\mathcal{M}_{d}$ for any $i$. Let $x$ be the input to the model at a single run, $\mathcal{M}_{t}(x)$ and $\mathcal{M}_{i}(x)$ are then the output probability distribution with input $x$.
Expected acceptance rate $\alpha=\mathbf{E}(1-\frac{1}{2}|\mathcal{M}_{t}(x)-\mathcal{M}_{i}(x)|)$ is the probability of draft generation by $\{\mathcal{M}_{i}\}$ being accepted by target model $\mathcal{M}_{t}$ (Leviathan et al., 2023).
Cost coefficient $c_{i}$ is the ratio of time for a single run of draft model $\mathcal{M}_{i}$ over target model $\mathcal{M}_{t}$.
Expected walltime improvement factor (EWIF) is the expected time improvement of an algorithm based on. We provide the formula for EWIF in an extended setting when multiple draft models are involved.
Theorem 4.1.
The expected walltime improvement factor (EWIF) by Algorithm 1 is $T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\sum_{i=0}^{k}\prod_{j=1}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$.
Proof.
Since every token will only be considered by the target model if all tokens before are accepted, a simple recursion leads to the probability of acceptance of the $i$-th token being $\prod_{j=1}^{i-1}\alpha_{j}$, and the $\prod_{j=1}^{0}\alpha_{j}=1$ representing the token generated by the target model. In the right-hand side, the numerator is the expected number of tokens generated, and the denominator is the time costs for each step of such a system, so the right-hand side is the expected efficiency improvement. ∎
We can immediately achieve a formula for speculative decoding by requiring all the draft models to be the same.
Corollary 4.2.
When $q_{1}=q_{2}=...=q_{k}$, let $\alpha_{i}=\alpha$ and $c_{i}=c$ for any $i$, the EWIF is $\frac{1-\alpha^{k+1}}{(1-\alpha)(c\cdot k+1)}$.
Furthermore, theorem 4.1 can be used to analyze the importance of the tokens in the drafting step.
Corollary 4.3.
The probability of $i$-th token being accepted is $\prod_{j=1}^{i}\alpha_{j}$. The derivative of EWIF with respect to $\alpha_{l}$ is $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{l}}=\frac{\sum_{i=l}^{k}\prod_{j=1,j\neq l}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$. Specifically, $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{1}}=\frac{\sum_{i=1}^{k}\prod_{j=2}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$ and $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{k}}=\frac{\prod_{j=1}^{k-1}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$.
Using the information provided by Leviathan et al. (2023), we calculat a simulated EWIF under the assumption that the event of acceptance by the target model is a Bernoulli trial. The results, shown in Table 1, indicate that speculative sampling with a horizontal cascade achieved better EWIF than vanilla speculative sampling under this assumption.
4.2 Analysis of Model Interpolation
The setting of cascade drafting can be simply extended using model interpolation that involves choosing between two models using randomness. Given two draft systems $d_{1}$ and $d_{2}$ (They can be two models generating autoregressive or two speculative decoding systems as used in vertical cascade) and a probability function $f$, an interpolation $d_{inter}$ is defined as $f(i)d_{1}+(1-f(i))d_{2}$. In other words, with probability $f(i)$, $d_{inter}$ is $d_{1}$, and with probability $1-f(i)$, $d_{inter}$ is $d_{2}$. As a result, $\alpha_{i}=f(i)\alpha_{d1}+(1-f(i))\alpha_{d2}=\alpha_{d2}+f(i)(\alpha_{d1}-\alpha_{d2})$ and $c_{i}=f(i)c_{d1}+(1-f(i))c_{d2}=c_{2}+f(i)(c_{d1}-c_{d2})$. An interpolation system can provide a smooth transition between two systems; however, in theorem 4.4, we prove that the optimal can always be reached without interpolation.
Theorem 4.4.
Given a horizontal cascade system with the setting mentioned above. There exists a probability function $f$ that optimizes the system such that $f$ satisfies $f(i)\in\{0,1\}$ for any $i$. This means each of the tokens will be generated by either $d_{1}$ or $d_{2}$, and the optimal system is deterministic.
Proof of Theorem 4.4 is included in appendix B.
Dataset $\mathcal{M}_{d1}$ $\mathcal{M}_{d2}$ $k_{1}$ $k_{2}$ EWIF
CNNDM small - 9 - 2.65
CNNDM base - 8 - 2.96
CNNDM base small 5 3 3.03
ENDE small - 12 - 3.61
ENDE base - 11 - 3.75
ENDE base small 5 8 3.93
Table 1: Simulated EWIF under the assumption that the acceptance distribution is a Bernoulli distribution. BASE and SMALL refer to FLAN-T5-base and FLAN-T5-small. In the simulation, speculative sampling with horizontal cascade exceeded the performance of the vanilla speculative decoding on both CNN Dailymail (Nallapati et al., 2016) and WMT EnDe (Bojar et al., 2018) datasets.
5 Evaluation
Previous works on speculative decoding and related methods relied on walltime as an evaluation method. However, there are standardization and legitimacy concerns related to walltime. Leviathan et al. (2023) mentioned that cost coefficiency, the ratio between the time for a single run of draft model and target model will vary from system to system which can result in significant variation in wall time improvement across different systems, making comparisons across different research work difficult.
In addition, a recent analysis suggests GPU speed can vary for the same GPU model with one being 1.5x faster than another GPU of the same model (Sinha et al., 2022), and even for the same GPU, the performance variation between different runs is unknown; therefore, theoretical walltime improvement up to 1.5x can be undermined by GPU variation, and methods without any improvement can be measured with an improvement up to 1.5x by running on a different GPU. The properties of such lack of standardization and stability make walltime unfit for measuring algorithmic improvement. To resolve this issue, we design a new metric to measure speculative decoding and its related algorithms which can reduce the impact of varying/unstable GPU performance.
Our proposed method, standardized walltime improvement (SWI), calculates the GPU times of the models, assuming that each run of a language model costs the same amount of time, an assumption made when inventing the speculative sampling algorithm (Chen et al., 2023). Given a speculative decoding system $S$ containing $k$ draft models $\mathcal{M}_{d1},...,\mathcal{M}_{dk}$, target model $\mathcal{M}_{t}$ solving a task $T$, and let $count(S,m,T)$ be the count of runs of model $m$ during the execution of task $T$ with system $S$, let $c_{i}$ be the expected cost coefficient being the time ratio between a single run of $\mathcal{M}_{i}$ and $\mathcal{M}_{t}$, and let $n$ be the total of tokens generated, the standardized walltime improvement is defined as
$\begin{split}SWI&(\mathcal{M}_{d1},...,\mathcal{M}_{dk},\mathcal{M}_{t},T)=\\ &\frac{n}{count(S,\mathcal{M}_{t},T)+\sum_{i=1}^{n}count(S,\mathcal{M}_{di},T)c_{i}},\end{split}$
which is simply the sum of GPU time by all models when $c_{i}$ is measured during a local experiment. The benefit of the $SWI$ evaluation is that $count(S,\mathcal{M}_{t},T)$ and $count(S,\mathcal{M}_{di},T)$ will be the same across different systems, so by using the same $c_{i}$ across different experiments the GPU performance variation will be eliminated; therefore, it provides a standardized, reproducible measurement.
During our experiment, we will use two sets of $\{c_{i}\}$, existing $\{c_{i}\}$ reported by Leviathan et al. (2023) and an heuristic $\{c_{i}\}$ being the ratio of number of tunable parameters between $\mathcal{M}_{di}$ and $\mathcal{M}_{t}$.
Dataset Algorithm $\{\mathcal{M}_{di}\}$ SWI (MS) $k_{11}$ $k_{21}$ $k_{12}$ $l$
GSM8K Autoregressive - 1 - - - -
GSM8K S. Decoding base $3.38$ 10 - - -
GSM8K S. Decoding small $3.06$ 11 - - -
GSM8K C.S. Drafting base, mag $3.62$ 9 - - -
GSM8K C.S. Drafting small, mag $3.16$ 11 - - -
GSM8K C.S. Drafting base, small, mag 3.75 7 1 10 1.5
MMLU Autoregressive - 1 - - - -
MMLU S. Decoding base $3.97$ 13 - - -
MMLU S. Decoding small $4.12$ 19 - - -
MMLU C.S. Drafting base, mag $4.50$ 12 - - -
MMLU C.S. Drafting small, mag $4.36$ 17 - - -
MMLU C.S. Drafting base, small, mag 4.83 5 1 19 2
Dataset Algorithm $\{\mathcal{M}_{di}\}$ SWI (PW) $k_{11}$ $k_{21}$ $k_{12}$ $l$
GSM8K Autoregressive - 1 - - - -
GSM8K S. Decoding base $2.99$ 8 - - -
GSM8K S. Decoding small $2.76$ 8 - - -
GSM8K C.S. Drafting base, mag $3.22$ 8 - - -
GSM8K C.S. Drafting small, mag $2.90$ 7 - - -
GSM8K C.S. Drafting base, small, mag 3.29 5 1 9 1.5
MMLU Autoregressive - 1 - - - -
MMLU S. Decoding base $3.42$ 10 - - -
MMLU S. Decoding small $3.51$ 11 - - -
MMLU C.S. Drafting base, mag $4.07$ 9 - - -
MMLU C.S. Drafting small, mag $3.97$ 12 - - -
MMLU C.S. Drafting base, small, mag 4.23 5 1 13 1.5
Table 2: The experiment results from datasets GSM8K and MMLU. BASE and SMALL refer to FLAN-T5-base and FLAN-T5-small. SWI (MS) is the standardized walltime improvement with the assumption that the latency of each run of a model is its number of parameters (model size). SWI (PW) is the SWI with the assumption that the latency of each run of a model is the time cost data reported from previous work (Leviathan et al., 2023). $k_{11}$, $k_{21}$, $k_{12}$, $l$ are the hyperparameters. $k_{1i}$ represents the step limitation target model and the draft models, $k_{21}$ is the step limitations between the first and second draft model, and $l$ is leniency as shown in algorithm 1. For speculative decoding, the $k_{11}$ is simply the $k$.
6 Experiment
6.1 Setup
We chose our target and draft models from the FLAN-T5 (Chung et al., 2022) family for our experiment, as there is a large variation in model sizes within the FLAN-T5 family (ranging from 77 million to 11 billion parameters). To align our experiment with current common usage, we do not perform fine-tuning, and the generation is conducted in a zero-shot manner. We use FLAN-T5-xxl as our target model, FLAN-T5-base and FLAN-T5-small as our reviewing draft models, and the Max-Gram algorithm as the generating draft model.
To ensure our experiments are fully reproducible, we perform sampling at temperature 00, which is essentially greedy decoding. And we do not observe any significant difference between sampling with temperature $1$ and greedy decoding in previous speculative decoding experiments (Leviathan et al., 2023).
Cascade Speculative Drafting Inference To reduce the number of hyperparameters to tune, we use MaG to generate 10 tokens at once, as it is rare for more than 10 tokens to be accepted. We do not use lenience when the reviewer is $\mathcal{M}_{t}$ to ensure the output distribution remains does not change. Additionally, we eliminate leniency between $\mathcal{M}_{d2}$ and MaG because the performance gap between them is large.
With these constraints, we are left with at most four hyperparameters: $k_{11}$, $k_{21}$, $k_{12}$, and $l$. For the CS Drafting step where the target model is the reviewer, $k_{11}$ and $k_{12}$ are used. $k_{21}$ and $l$ are used in the step where $\mathcal{M}_{d1}$ is the reviewer. However, there are still too many hyperparameters for us to exhaustively search for the optimal result; instead, we employ a halving grid search to find a set of hyperparameters.
Baseline We consider both speculative decoding and vanilla autoregressive generation as our baselines. For speculative decoding, we opt for FLAN-T5-small and FLAN-T5-base as the draft models, as they perform better than FLAN-T5-large, according to Leviathan et al. (2023) and our preliminary experiments. We test a wide range of $k$-values, from 2 to 30, to ensure that we find the best result for each model serving as our draft model to establish our baseline.
Dataset We chose two commonly used datasets for our experiments. For both datasets, we conducted experiments in a zero-shot chain-of-thought setup (Kojima et al., 2022; Wei et al., 2023):
  • GSM8K (Cobbe et al., 2021) is a dataset comprising 8,500 high-quality, linguistically diverse, grade-school math word problems. It focuses on multi-step reasoning with problems that are typically solvable using basic arithmetic in 2-8 steps.
  • MMLU (Hendrycks et al., 2021), or Massive Multitask Language Understanding, is a benchmark for testing how well large language models grasp knowledge. It encompasses 57 diverse subjects, ranging from elementary science to advanced law.
6.2 Experimental Results
Table 2 presents the experimental results. In two settings of SWI, Cascade Speculative Decoding has outperformed the speculative decoding algorithm. For GSM8K, Cascade Speculative Decoding achieved a maximum additional speedup of 37% over the fastest speculative algorithm; for MMLU, the maximum additional speedup improvement over speculative decoding is 72%.
Additionally, we emphasize the effectiveness of the MaG algorithm. When comparing CS Drafting with one neural model and MaG against the fastest speculative decoding setup, we found that CS Drafting with one neural model gained up to a 56% speedup on MMLU and a 24% speedup on GSM8K. Notably, the MaG algorithm only involves a bigram model with as many parameters as the size of the tokenizer, making its memory cost negligible. Thus, the speedup gained using CS Drafting with one neural model involves no additional deployment overhead while reducing both latency and computational cost, making it a superior choice over speculative decoding.
When comparing CS Drafting with two neural models to CS Drafting with one neural model, the former requires additional memory utilization. Despite the latency reduction, CS Drafting with two neural models can potentially incur more computational cost. However, during our experiments, since the draft models had only 2% of the number of parameters of the target model, the memory cost did not outweigh the reduced latency. Therefore, the CS Drafting algorithm with two neural models still uses less computational resources, as it completes each generation round faster.
Another observation is that despite FLAN-T5-small sometimes outperforming FLAN-T5-base as a draft model for speculative decoding, in CS Drafting with the aid of MaG, FLAN-T5-base consistently outperforms FLAN-T5-small. This raises a topic for future study: determining the ideal size of draft models for CS Drafting.
7 Related Work
7.1 Efficienct Methods for Language Model Inference
In the era of large language models, the efficiency during inference becomes a key to model service. To reduce the model inference cost and speed up, several efficient methods have been proposed, including pruning, knowledge distillation and quantization (Treviso et al., 2023). Model pruning takes structured (Xia et al., 2022; Voita et al., 2019) or unstructured (Guo et al., 2021; Chen et al., 2020) methods to remove the redundant model parameters to reduce the storage memory and increase inference speed. Knowledge distillation takes the approach of transferring knowledge from a superior teacher model to a smaller student model (Hinton et al., 2015; Gou et al., 2021). Quantization maps high-precision data representations (e.g. 32 bits) into low-precision ones (e.g. 8 bits) to reduce memory consumption (Bhandare et al., 2019; Schaefer et al., 2023).
7.2 Speculative Decoding
Recent works improved the speed of speculative decoding by experimenting with task-specific knowledge distillation (Zhou et al., 2023), or using approaches similar to speculative decoding to achieve a tradeoff between quality and speed (Leviathan et al., 2023; Kim et al., 2023). However, the reduced quality for speed can be undesirable, and task-specific training might not be applicable to many real-world NLP applications such as chatbots; overall, we found no generic improvement for speculative decoding in the current literature.
Speculative decoding (Chen et al., 2023; Leviathan et al., 2023) is proposed to improve inference speed by using a draft model to perform auto-regressive decoding and a target model to check and replace the generated tokens. A key to further speed up is to increase the acceptance rate in speculative decoding, which can be done by aligning the distribution between the draft model and the target model. Some recent work has proposed to improve the performance of speculative decoding in a domain/task-specific setting. Zhou et al. (2023) use generalized knowledge distillation to finetune the draft model and achieve 10%-45% speedup compared to draft models without knowledge distillation. Liu et al. (2023) propose Online Speculative Decoding to continually update multiple draft models on observed user query data, and the draft models align with the query distribution in real-time. However, both approaches require additional task-specific training making it difficult to apply the methods in a general setting. On the other hand, He et al. (2023) proposed constructing a datastore from a domain-specific corpus to serve as a draft model which can be used when a draft model is not available.
8 Conclusion
In our work, we propose a novel algorithm, CS Drafting, which involves two cascades: the vertical cascade and the horizontal cascade. The vertical cascade eliminates the necessity of autoregressive generation from a neural language model, while the horizontal cascade effectively allocates the cost of drafting tokens at different positions. Our experiments demonstrate that CS Drafting achieves up to a 72 percent additional speedup over speculative decoding without any additional training, while maintaining the same generation quality as the target model.
Appendix A Max-Gram Implementation
Listing 1: Max-Gram Algorithm
1def torch_index(t, value):
2    return (t == value).nonzero(as_tuple=True)[0][0]
3
4
5def max_gram(input_ids, encoder_ids, n=1):
6    matches = (encoder_ids[0] == input_ids[0, -1]).int()
7    if matches.sum() < 1:
8        return None
9    for i in range(2, input_ids.shape[-1] + 1):
10        new_matches = (encoder_ids[0, :(-1 * (i - 1))] == input_ids[0, -1 * i]).int()
11        combined_matches = (2 - new_matches == matches[1:]).int()
12        if combined_matches.sum() < 1:
13            index = torch_index(torch.cat(
14                (
15                    torch.tensor([0] * (i - 1), device=torch.device(encoder_ids.device)),
16                    matches
17                ),
18                dim=-1
19            ), 1)
20            return encoder_ids[:, index:index + n]
21        else:
22            matches = combined_matches
23    index = torch_index(torch.cat((
24        torch.tensor([0] * (encoder_ids.shape[-1] - matches.shape[-1])), matches), dim=-1
25        ), 1)
26    return encoder_ids[:, index+1:index + n+1]
Appendix B Proof and Analysis of Theorem 4.4
We first prove a lemma that helps the proof of theorem 4.4
Lemma B.1.
$\forall a,b,c,d\in\mathbb{R}_{>0},\gamma\in(0,1)$
$\frac{\gamma a+(1-\gamma)b}{\gamma c+(1-\gamma)d}\leq\max(\frac{a}{c},\frac{b}{d})$
Proof.
Without loss of generality, assume $\frac{a}{c}\geq\frac{b}{d}$,
$\frac{a}{c}\geq\frac{\gamma a+(1-\gamma)b}{\gamma c+(1-\gamma)d}$
$\Longleftrightarrow a(\gamma c+(1-\gamma)d)\geq c(\gamma a+(1-\gamma)b)$
$\Longleftrightarrow(1-\gamma)ad\geq(1-\gamma)cb$
$\Longleftrightarrow\frac{a}{c}\geq\frac{b}{d}$
Proof of Theorem 4.4
Proof.
Let $\alpha_{d_{1}},\alpha_{d2}$ and $c_{d_{1}},c_{d2}$ be the acceptence rate and cost ratio for model $d_{1},d_{2}$ respectively. Without loss of generality, we assume $\alpha_{d_{1}}>\alpha_{d2}$ and $c_{d_{1}}>c_{d2}$ Then the system satisfies $\alpha_{i}=f(i)\alpha_{d_{1}}+(1-f(i))\alpha_{d_{2}}$ and $c_{i}=f(i)\alpha_{d_{1}}+(1-f(i))c_{d_{2}}$ We prove using contradiction. Assuming all optimal systems need at least one interpolation of models at some positions. Let an optimal system of $f$ given $d_{1},d_{2}$, Without loss of generality, assume $r$ to be the first indices such that $f(r)$ not being an integer and the system $f$ to be the one with largest $r$ among all optimal systems.
$T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\alpha_{r}(1+\sum_{i=r+1}^{k}\prod_{j=r+1}^{i}\alpha_{j})+\sum_{i=0}^{r-1}\prod_{j=1}^{i}\alpha_{j}}{1+c_{r}+\sum_{i=1,i\neq r}^{k}c_{i}}$
$=\frac{(f(r)\alpha_{d_{1}}+(1-f(r))\alpha_{d_{2}})(1+\sum_{i=r+1}^{k}\prod_{j=r+1}^{i}\alpha_{j})+\sum_{i=0}^{r-1}\prod_{j=1}^{i}\alpha_{j}}{1+f(r)c_{d_{1}}+(1-f(r))c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i}}$
Let $A_{1}=1+\sum_{i=r+1}^{k}\prod_{j=r+1}^{i}\alpha_{j},A_{2}=\sum_{i=0}^{r-1}\prod_{j=1}^{i}\alpha_{j}$
$=\frac{f(r)(\alpha_{d_{1}}A_{1}+A_{2})+(1-f(r))(\alpha_{d_{2}}A_{1}+A_{2})}{f(r)(1+c_{d_{1}}+\sum_{i=1,i\neq r}^{k}c_{i})+(1-f(r))(1+c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i})}$
$\leq\max{(\frac{\alpha_{d_{1}}A_{1}+A_{2}}{1+c_{d_{1}}+\sum_{i=1,i\neq r}^{k}c_{i}},\frac{\alpha_{d_{2}}A_{1}+A_{2}}{1+c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i}})}$
This means that $f(r)=0$ or $f(r)=1$ will result in a time improvement factor that is no less than that of a non-integer $f$, indicating that replacing $f(r)$ with an integer will still yield an optimal system. This contradicts the assumption that $f$ is the system with the largest $r$. As a result, there exists an optimal system that does not require interpolation. ∎
Figure 3: Acceptance rate of draft tokens by their positions on MMLU datasets, evaluated on FLAN-T5-small, base and large models.